OpenStudy (anonymous):

The result of dividing a number of two digits by the number with digits reversed is 5/6. If the difference of digits is 1, find the number.

OpenStudy (anonymous):

have u tried it yet?

OpenStudy (anonymous):

Yes, 10x+y/10y+x=5/6 Which implies 60x+6y=50y+5x 55x-44y=0 How to proceed??

OpenStudy (anonymous):

the difference of digits is 1 so x=y+1 or y=x+1

OpenStudy (anonymous):

so u have a system of equations\[5x-4y=0\]\[x-y=1\]or\[5x-4y=0\]\[y-x=1\]

OpenStudy (anonymous):

makes sense???

OpenStudy (anonymous):

How come 5x-4y=0? The second equation makes sense.

OpenStudy (anonymous):

well 55x-44y=0 divide by 11 :)

OpenStudy (anonymous):

lol! Thanks a lot !! :D

OpenStudy (anonymous):

:D yw :)